# Vision Mamba in Remote Sensing: A Comprehensive Survey of Techniques, Applications and Outlook

### v1

@@ IV-B2 Scan Sampling
Figure 6: The four scam sampling for scan strategy.
@@ IV-B3 Scan Directions
Similarly, certain preprocessing methods, such as SaT, can directly transform 2D feature maps into 1D sequences, bypassing some subsequent scan steps. For 1D sequences, scanning options include two starting tokens, namely the first token ($\leftarrow$) and the last token ($\rightarrow$). In addition, the initial direction is from left to right if starting token is the first, and is from right to left if starting token is the last, which is recorded as $\rightarrow$.
@@ IV-B4 Scan Patterns
After completing previous steps, the subsequent step involves employing suitable scan patterns to systematically convert the 2D feature map into 1D sequences. Various scan patterns provide unique benefits tailored to specific applications. In remote sensing, 10 primary scan patterns are commonly used, which are visually illustrated in Fig. 8.
Figure 8: The ten scam pattern for scan strategy.
@@ IV-C Multi-Modal and Bi-Temporal Feature Interaction via Mamba
Numerous studies [81, 24, 123, 109, 126, 35, 36, 25, 82, 111, 80, 151, 117] have leveraged multimodal and bi-temporal data, such as conventional RGB images, LiDAR point clouds and infrared images, to enhance performance. The bi-temporal data refers two images captured at different time in the same location, normally for the Change Detection (CD) task.
@@ IV-C3 Scan Strategy-based Feature Post-Processing Methods
Similarly, following post-processing methods of the scan strategy framework in Section IV-B5, a total of four methods [123, 126, 151, 117, 24] fall under the category of scan strategy’s feature post-processing-based methods. Unlike preprocessing-based methods, these approaches are applied after obtaining 1D sequences.
@@ V-A1 Hybrid Architecture with CNNs
The integration of CNNs with Mamba aims to complement the global modeling capability of Mamba with local feature extraction capability of CNN. Various approaches [28, 29, 81, 89, 87, 24, 124, 125, 34, 154, 127, 128, 121, 88, 129, 25, 91, 130, 79, 156, 26, 120, 142, 147, 93, 143, 95, 119, 82, 97, 98, 178, 155, 152, 100, 175, 84, 145, 177, 146, 101, 102, 103, 153, 149, 107, 80, 133, 134, 150, 104, 159, 116, 162, 108, 148, 136, 137, 138, 106, 139, 161, 158, 122] have been proposed to achieve this hybridization, typically falling into sequential or parallel configurations—both within each stackable building block and across the overall architectural framework.
@@ V-C Learning Paradigm
This subsection provides an overview of various learning paradigms applied in remote sensing, with emphasis on unsupervised, self-supervised and prompt learning paradigms. Except for these paradigms, the self-supervised learning is used by the rest of the work.
@@ V-D1 Fast Fourier Transform
FreMamba [32] is the first to merge Mamba with Fourier transformation within its basic unit, adding an FFT-based branch sequentially to both the Mamba-based and CNN-based blocks. This design facilitates the exploration of spatial and frequency correlations. Similarly, FMambaIR [131] applies an attention mechanism separately to the amplitude and phase obtained via FFT, alongside an additional Mamba branch, thereby enhancing the extraction of global degradation features and overall global information perception. MambaFormerSR [130] introduces an FFT-based attention mechanism into the conventional Feed-Forward Network (FFN). By multiplying the FFT-enhanced branch (processed through CNNs and FFT) with a branch solely processed by CNNs, the network is guided to emphasize degradation-sensitive components during image restoration. In TransMamba [163], a dual-branch architecture comprised Transformer and Mamba modules is proposed. Within the Transformer branch, a self-attention module applies FFT to reallocate features into distinct frequency bands. Learnable attention weights then adaptively suppress low-frequency rain streaks while amplifying high-frequency textures, and a Spectral Enhanced Feed-Forward (SEFF) module further refines features with frequency-specific filters and dilated convolutions.
@@ VI Downstream Applications
Method Rate OA AA Kappa
@@
TABLE III: The classification benchmark for the Pavia University Dataset [202]. Rate represents the percentage of the training data to the total data; OA represents overall classification accuracy; AA represents average classification accuracy; Kappa represents the kappa coefficient.
@@
TABLE IV: The segmentation benchmark for the ISPRS Vaihingen dataset [213]. mF1 represents mean F1 score; mIOU represents mean intersection over union; OA represents overall accuracy.
@@
TABLE V: The change detection benchmark for WHU-CD dataset [224].
@@
TABLE VI: The object detection benchmark for DroneVehicle [225] and VisDrone [226] datasets. mAP represents mean average precision; mAP0.5 represents the mAP at an IoU threshold of 50%. $\mathcal{M}$ and $\mathcal{Y}$ represent Mamba and YOLO for short. represents use of multimodal data, RGB and thermal infrared images.
@@
TABLE VII: The super-resolution benchmark for Chikusei [239] and AID [240] datasets (scale factor is 4 $\times$). PSNR represents peak signal-to-noise ratio and SSIM represents structural similarity index. The data on Chikusei [239] are preprocessed differently so that the performances of HSRMamba [122] and MambaIR [249] are better.
@@
TABLE VIII: The image restoration benchmark for SateHaze1k [258] and UAV-Rain [259].
In this section, we introduce six benchmarks involving models based on CNN, Transformer, and Mamba architectures, across various downstream tasks. To provide a more comprehensive comparison and demonstration, we have selected tasks that include both high-level vision tasks (i.e., image classification (Tab. III), segmentation (Tab. IV), change detection (Tab. V, object detection (Tab. VI) and low-level vision tasks (i.e., super-resolution (Tab. VII), image restoration (Tab. VIII)).
Results on Image Classification Task: Tab. III shows the classification performance among different models on the Pavia University Dataset [202]. As we can see, although the Rate parameters set by each method vary, when the same Rate (5.00%) is adopted, the IGroupSS-Mamba [167] achieves optimal performance in parameters such as OA, AA, Kappa, and surpasses all other CNN-based and Transformer-based models.
Results on Image Segmentation Task: Tab. IV presents the comparison performance among different models on the ISPRS Vaihingen dataset [213]. Overall, Transformer-based methods tend to outperform CNN-based methods, but Mamba-based methods generally surpass Transformer-based methods in most cases. Specifically, PPMamba [142] achieves the best performance on the mF1 and mIOU metrics, while [128] excels in the OA metric among all methods reporting this metric.
Results on Image Change Detection Task: Tab. V shows the performance comparisons among different models on the WHU-CD dataset [224]. Similarly, Mamba-based models achieve better performance than CNN-based and Transformer-based models. Specifically, DC-Mamba [124] delivers the best performance in the F1 metric, TTMGNet [84] excels in the IoU metric, and ChangeFormer [221] achieves the best performance in the OA metric. For more comprehensive comparison, it is recommended that the author refer to relevant survey paper [174].
Results on Image Object Detection Task: Tab. VI compares the performance of object detection tasks on the DroneVehicle [225] and VisDrone [226] datasets. Specifically, the Mamba-based model, ES-HS-FPN [102], outperforms both CNN-based and Transformer-based models, achieving the highest mAP value.
Results on Image Super-Resolution Task: Table VII presents a comparison of image super-resolution task performance across the Chikusei [239] and AID [240] datasets. In particular, FreMamba [32] records the highest values in both PSNR and SSIM metrics on the AID dataset.
Results on Image Restoration Task: Tab. VIII presents a performance comparison of image restoration tasks on the SateHaze1k [258] and UAV-Rain [259] datasets. Notably, RSDehamba [31] sets a new state-of-the-art benchmark for the tasks on the SateHaze1k dataset in both PSNR and SSIM metrics.
@@ VII-A Causality
The S6 block in Mamba operates as a causal system, wherein predictions for each token are generated based solely on the current input and a hidden state summarizing preceding information. This architecture is inherently optimized for one-dimensional sequential data rather than two-dimensional image. Although patch embedding techniques and flattening allow to transform 2D data into 1D sequences, this process inevitably results in the loss of spatial information due to the causal nature of the system, thereby compromising performance.
@@ VII-C Multi-Modal Interaction via Mamba
For data rich in spectral information, such as hyperspectral image, an intriguing approach is to consider these data as 3D data. Within the Mamba framework, only SSUMamba [153] currently leverages a 3D scan strategy to process such data. Further exploration of 3D scanning techniques to capture spatial-spectral relationships represents an unexplored but promising direction in remote sensing research.
@@ VII-E Mamba-based Foundation Models in Remote Sensing
Nonetheless, scaling Mamba-based models to larger sizes may introduce stability issues [40]. Although several backbone networks [64, 75, 76, 21] have proposed methods to enhance stability in large-scale implementations, these solutions have not yet been fully validated for foundation models in remote sensing. Addressing the stability challenges associated with large Mamba-based models is therefore a crucial area for future research.
@@ VII-G Adaptations to Downstream Tasks
While Mamba-based models have demonstrated promise in several remote sensing applications, their full potential remains underexplored, particularly in unconventional or emerging downstream tasks. The architecture’s advantages, such as linear scaling and efficient long-range dependency modeling, could effectively address unique challenges in less-studied tasks, such as some application in agriculture [261], in forest [262] and in ecological restoration [263] and in less-explored data, such as hyperspectral time-series analysis [264] that is characterized by both high spectral resolution and temporal continuity. Moreover, the potential of vision-language (CV + NLP), for remote sensing remains underexplored. A Mamba-based multimodal framework could enable novel applications such as content-guided image retrieval [265] in remote sensing data , automated visual question answering (VQA) [266] for non-expert users and image captioning [267] of remote sensing imagery. In addition, given the recent success of Large Language Models (LLMs), combining Vision Mamba with LLMs represents a promising direction, as it leverages the strong reasoning capabilities of LLMs. Beyond the tasks discussed in Section VI, further investigation into applications traditionally addressed by CNN and Transformer models represents a significant and promising research direction.

### v2

@@ IV-B2 Scan Sampling
Figure 7: The four scan sampling for the scan strategy.
@@ IV-B3 Scan Directions
Similarly, certain preprocessing methods, such as SaT, can directly transform 2D feature maps into 1D sequences, bypassing some subsequent scan steps. For 1D sequences, scanning options include two starting tokens, namely the first token ($\leftarrow$) and the last token ($\rightarrow$). In addition, the initial direction is from left to right if the starting token is the first, and is from right to left if the starting token is the last, which is recorded as $\rightarrow$.
@@ IV-B4 Scan Patterns
After completing the previous steps, the subsequent step involves employing suitable scan patterns to systematically convert the 2D feature map into 1D sequences. Various scan patterns provide unique benefits tailored to specific applications. In remote sensing, 10 primary scan patterns are commonly used, which are visually illustrated in Fig. 9.
Figure 9: The ten scan patterns for the scan strategy.
@@ IV-C Multi-Modal and Bi-Temporal Feature Interaction via Mamba
Numerous studies [81, 24, 123, 109, 126, 35, 36, 25, 82, 111, 80, 151, 117] have leveraged multimodal and bi-temporal data, such as conventional RGB images, LiDAR point clouds and infrared images, to enhance performance. The bi-temporal data refers to two images captured at different time in the same location, normally for the Change Detection (CD) task.
@@ IV-C3 Scan Strategy-based Feature Post-Processing Methods
Similarly, following post-processing methods of the scan strategy framework in Section IV-B5, a total of four methods [123, 126, 151, 117, 24] fall under the category of the scan strategy’s feature post-processing-based methods. Unlike preprocessing-based methods, these approaches are applied after obtaining 1D sequences.
@@ V-A1 Hybrid Architecture with CNNs
The integration of CNNs with Mamba aims to complement the global modeling capability of Mamba with the local feature extraction capability of CNNs. Various approaches [28, 29, 81, 89, 87, 24, 124, 125, 34, 154, 127, 128, 121, 88, 129, 25, 91, 130, 79, 156, 26, 120, 142, 147, 93, 143, 95, 119, 82, 97, 98, 178, 155, 152, 100, 175, 84, 145, 177, 146, 101, 102, 103, 153, 149, 107, 80, 133, 134, 150, 104, 159, 116, 162, 108, 148, 136, 137, 138, 106, 139, 161, 158, 122] have been proposed to achieve this hybridization, typically falling into sequential or parallel configurations—both within each stackable building block and across the overall architectural framework.
@@ V-C Learning Paradigm
This subsection provides an overview of various learning paradigms applied in remote sensing, with emphasis on unsupervised, self-supervised and prompt learning paradigms. Except for these paradigms, self-supervised learning is used for the rest of the work.
@@ V-D1 Fast Fourier Transform
FreMamba [32] is the first to merge Mamba with the Fourier transformation within its basic unit, adding an FFT-based branch sequentially to both the Mamba-based and CNN-based blocks. This design facilitates the exploration of spatial and frequency correlations. Similarly, FMambaIR [131] applies an attention mechanism separately to the amplitude and phase obtained via FFT, alongside an additional Mamba branch, thereby enhancing the extraction of global degradation features and overall global information perception. MambaFormerSR [130] introduces an FFT-based attention mechanism into the conventional Feed-Forward Network (FFN). By multiplying the FFT-enhanced branch (processed through CNNs and FFT) with a branch solely processed by CNNs, the network is guided to emphasize degradation-sensitive components during image restoration. In TransMamba [163], a dual-branch architecture comprised Transformer and Mamba modules is proposed. Within the Transformer branch, a self-attention module applies FFT to reallocate features into distinct frequency bands. Learnable attention weights then adaptively suppress low-frequency rain streaks while amplifying high-frequency textures, and a Spectral Enhanced Feed-Forward (SEFF) module further refines features with frequency-specific filters and dilated convolutions.
@@ VI Downstream Applications
TABLE III: The overview of five tasks in remote sensing and their representative work, including classification, semantic segmentation, object detection, change detection, super-resolution and image restoration.
Task Representative Works
Classification
[36], [79], [81], [83], [91], [92], [94], [95], [96]
[99], [100], [101], [104], [105], [106], [109]
[110], [112], [113], [114], [116], [118], [119]
[120], [144], [152], [156], [157], [165], [167]
[177], [178], [179], [191], [192],
Semantic
Segmentation
[27], [28], [29], [82], [88], [90], [93], [108]
[121], [125], [126], [127], [128], [134], [140]
[141], [142], [145], [154], [155], [181],  [158]
Object
Detection
[24], [25], [26], [80], [102], [115], [136]
[148], [162], [193]
Change
Detection
[27], [87], [89], [84], [123], [124], [138], [146]
[147], [151], [175]
Super-
Resolution
[32], [97], [98], [122], [129], [130], [143], [168]
Image
Restoration
[31], [33], [34], [107], [131], [160], [132]
[150], [153], [161], [163], [164], [180]
TABLE IV: The classification benchmark for the Pavia University Dataset [194]. Rate represents the percentage of the training data to the total data; OA represents overall classification accuracy; AA represents average classification accuracy; Kappa represents the kappa coefficient.
Methods Rate OA AA Kappa
@@
TABLE V: The segmentation benchmark for the ISPRS Vaihingen dataset [206]. mF1 represents mean F1 score; mIOU represents mean intersection over union; OA represents overall accuracy.
@@
TABLE VI: The change detection benchmark for WHU-CD dataset [217].
@@
TABLE VII: The object detection benchmark for DroneVehicle [228] and VisDrone [229] datasets. mAP represents mean average precision; mAP0.5 represents the mAP at an IoU threshold of 50%. $\mathcal{M}$ and $\mathcal{Y}$ represent Mamba and YOLO for short. represents the use of multimodal data, RGB and thermal infrared images.
@@
TABLE VIII: The super-resolution benchmark for Chikusei [242] and AID [243] datasets (scale factor is 4 $\times$). PSNR represents peak signal-to-noise ratio and SSIM represents structural similarity index. The data on Chikusei [242] are preprocessed differently so that the performances of HSRMamba [122] and MambaIR [244] are better.
@@
TABLE IX: The image restoration benchmark for SateHaze1k [253] and UAV-Rain [254].
@@
In this section, we introduce six benchmarks involving models based on CNN, Transformer, and Mamba architectures, across various downstream tasks. Tab. III lists five main tasks in remote sensing and their representative works. To provide a more comprehensive comparison and demonstration, we have selected tasks that include both high-level vision tasks (i.e., image classification (Tab. IV), segmentation (Tab. V), change detection (Tab. VI, object detection (Tab. VII) and low-level vision tasks (i.e., super-resolution (Tab. VIII), image restoration (Tab. IX)).
Results on Image Classification Task: Tab. IV shows the classification performance among different models on the Pavia University Dataset [194]. As we can see, although the Rate parameters set by each method vary, when the same Rate (5.00%) is adopted, the IGroupSS-Mamba [167] achieves optimal performance in parameters such as OA, AA, Kappa, and surpasses all other CNN-based and Transformer-based models.
Results on Image Segmentation Task: Tab. V presents the comparison performance among different models on the ISPRS Vaihingen dataset [206]. Overall, Transformer-based methods tend to outperform CNN-based methods, but Mamba-based methods generally surpass Transformer-based methods in most cases. Specifically, PPMamba [142] achieves the best performance on the mF1 and mIOU metrics, while [128] excels in the OA metric among all methods reporting this metric.
Results on Image Change Detection Task: Tab. VI shows the performance comparisons among different models on the WHU-CD dataset [217]. Similarly, Mamba-based models achieve better performance than CNN-based and Transformer-based models. Specifically, DC-Mamba [124] delivers the best performance in the F1 metric, TTMGNet [84] excels in the IoU metric, and ChangeFormer [225] achieves the best performance in the OA metric. For more comprehensive comparison, it is recommended that the author refer to the relevant survey paper [174].
Results on Image Object Detection Task: Tab. VII compares the performance of object detection tasks on the DroneVehicle [228] and VisDrone [229] datasets. Specifically, the Mamba-based model, ES-HS-FPN [102], outperforms both CNN-based and Transformer-based models, achieving the highest mAP value.
Results on Image Super-Resolution Task: Table VIII presents a comparison of image super-resolution task performance across the Chikusei [242] and AID [243] datasets. In particular, FreMamba [32] records the highest values in both PSNR and SSIM metrics on the AID dataset.
Results on Image Restoration Task: Tab. IX presents a performance comparison of image restoration tasks on the SateHaze1k [253] and UAV-Rain [254] datasets. Notably, RSDehamba [31] sets a new state-of-the-art benchmark for the tasks on the SateHaze1k dataset in both PSNR and SSIM metrics.
@@ VII-A Causality
The S6 block in Mamba operates as a causal system [20], wherein predictions for each token are generated based solely on the current input and a hidden state summarizing preceding information. This architecture is inherently optimized for one-dimensional sequential data rather than two-dimensional images. Although patch embedding techniques and flattening allow to transform 2D data into 1D sequences, this process inevitably results in the loss of spatial information due to the causal nature of the system, thereby compromising performance [37, 39, 40, 1, 41, 38].
@@ VII-C Multi-Modal Interaction via Mamba
For data rich in spectral information, such as the hyperspectral image, an intriguing approach is to consider these data as 3D data. Within the Mamba framework, only SSUMamba [153] currently leverages a 3D scan strategy to process such data. Further exploration of 3D scanning techniques to capture spatial-spectral relationships represents an unexplored but promising direction in remote sensing research.
@@ VII-E Mamba-based Foundation Models in Remote Sensing
Nonetheless, scaling Mamba-based models to larger sizes may introduce stability issues [40]. Although several backbone networks [64, 75, 76, 21] have proposed methods to enhance stability in large-scale implementations, these solutions have not yet been fully validated for foundation models in remote sensing. Addressing the stability challenges associated with large Mamba-based models is thus a crucial area for future research.
@@ VII-G Adaptations to Downstream Tasks
While Mamba-based models have demonstrated promise in several remote sensing applications, their full potential remains underexplored, particularly in unconventional or emerging downstream tasks. The architecture’s advantages, such as linear scaling and efficient long-range dependency modeling, could effectively address unique challenges in less-studied tasks, such as some application in agriculture [264], in forest [265] and in ecological restoration [266] and in less-explored data, such as hyperspectral time-series analysis [267] that is characterized by both high spectral resolution and temporal continuity. Moreover, the potential of vision-language (CV + NLP) for remote sensing remains underexplored. A Mamba-based multimodal framework could enable novel applications such as content-guided image retrieval [268] in remote sensing data, automated visual question answering (VQA) [269] for non-expert users and image captioning [270] of remote sensing imagery. In addition, given the recent success of Large Language Models (LLMs), combining Vision Mamba with LLMs represents a promising direction, as it leverages the strong reasoning capabilities of LLMs. Beyond the tasks discussed in Section VI, further investigation into applications traditionally addressed by CNN and Transformer models represents a significant and promising research direction.